# Personal LLM Agents: Insights and Survey about the Capability, Efficiency and Security

### v1

@@ 4.2.2 Sensing Targets
  • Scene sensing is often readily perceptible but hold significant importance, leading to variations both in behavior and emphasis. For behavior instance, detecting a user in a library prompts the agent to adjust the phone to silent mode, while in a bar increasing the volume and activating vibration may be necessary. Similarly to emphasis, when a user is in a meeting room, the agent should focus more on tasks related to meeting content recording and work organization, whereas in a gym, emphasis should shift towards fitness plans and heart rate analysis. Previous work in scene awareness has employed various techniques, such as location-based approaches [148], audio or video analysis [149, 150], and sensor capabilities analyzing aspects like airflow through smartphone microphones to assess ventilation [126], or scene recognition achieved by analyzing macro photographs taken with the smartphone camera when placed near a surface [127].
  • Occasion perception is more elusive in perception, and their impacts are relatively discreet. Earlier studies have identified differences in behavior and emotion recognition tasks across countries [151] and regions [152]. The national, ethnic, religious, and cultural backgrounds implied by the current user and setting are crucial. Perceiving others and objects in the current environment is equally vital. For example, previous work detected social scenarios based on sensor data, analyzing the behavior of socially anxious individuals in different social settings [153]. Other research delved into analyzing drinking-related social scenes using multiple sensors, even predicting the size and gender composition of drinking groups [154]. Additionally, studies explored the relations between sensor data, dietary habits, and social settings, revealing a strong association between binge eating and social environments, making it predictable [155].
With the recent advancements in LLMs, there have also emerged some environment sensing methods. For example, forecasting pedestrian flow through the analysis of public events [156] and LLM-based environmental understanding using multiple sensors [157].
@@
  • Short-term user sensing encompasses various aspects, including users’ routine actions [158], or specialized activities such as toothbrushing effectiveness [159], user states such as working or resting [138, 135], user health conditions [160, 125, 161], as well as user emotions [162, 134] and stress levels [163]. Short-term sensing typically involve rapidly changing and shallow-level state information. Efficiently capturing such information can significantly enhance the context awareness of Personal LLM Agents.
  • Long-term user sensing mainly focus on the analysis of users’ profile and personality. Vairous approaches have been proposed to understand users’ work, study, and daily life. For instance, a study utilized sensor data from new smartphones to detect the prolonged psychological states of freshmen [164]. Another study demonstrated the capability to predict learning performance and social activities based on perception data [165]. Gao et al. [166] delve into the techniques to predict personality based on the intensity of physical activities. There is also research examining the relationship between sensor data and user career advancement [167], as well as a study that predicts user life satisfaction [168]. Furthermore, specific states of users have been a focus, including studies on the perception of mental illnesses [169], such as one that predicts and analyzes schizophrenia [170], and another that detects habits like smoking [171]. Long-term sensing involve deep and abstract information, containing the profound logic behind user behavior. These pieces of information are often more subtle, making perception and maintenance challenging. However, they constitute an essential aspect for advanced personal agents.
In the realm of user sensing, there are also several LLM-based initiatives, such as employing LLM for recommendation tasks [172, 173], sentiment analysis with LLM [174], and the development of a personal doctor equipped with inquiry and perception capabilities [175].
@@ 5 Efficiency
Anagnostidis et al. [233], Zhang et al. [234], etc. , leaf, text width=27em ] ] [Kernel Optimization, fill=blue!10 [ FlashAttention [235, 236], FlashDecoding++ [237], etc. , leaf, text width=24em ] ] [Speculative Decoding, fill=blue!10 [ Chen et al. [238], Leviathan et al. [239] , etc. , leaf, text width=20em ] ] ] [Memory Reduction (§5.1.3), fill=blue!10 [KV Quantization, fill=blue!10 [ ZeroQuant [219], SmoothQuant [220], etc. , leaf, text width=18em ] ] [KV Pruning, fill=blue!10 [ Anagnostidis et al. [233], Zhang et al. [234], etc. , leaf, text width=22em ] ] [Offloading, fill=blue!10 [ FlexGen [240], etc. , leaf, text width=9em ] ] ] [Energy Optimization (§5.1.4), fill=blue!10 [Software Approaches, fill=blue!10 [ Same above, leaf, text width=8em ] ] [Hardware Approaches, fill=blue!10 [ NPU [241], TPU [242], FPGA [243], etc. , leaf, text width=18em ] ] ] ] [Efficient
@@
Manipulation (§5.3), fill=blue!10 [Searching Efficiency (§5.3.1), fill=blue!10 [Search Mechanism Design, fill=blue!10 [ Search Plan: rule-based [253, 254], cost-based [255, 256], etc. , leaf, text width=26em ] [ Metadata Filtering: AnalyticDB-V [255], HQANN [257], etc. , leaf, text width=26em ] ] [Search Process Execution, fill=blue!10 [ OPENMP: Faiss [258], Milvus [256], etc. , leaf, text width=18em ] [ SIMD: Faiss [258], Milvus [256], Quicker ADC [259], etc. , leaf, text width=25em ] [ GPU: Faiss [258], Milvus [256], etc. , leaf, text width=16em ] [ Distributed: Vald [260], Qdrant [253], etc. , leaf, text width=18em ] ] ] [Index Optimization (§5.3.2), fill=blue!10 [Indexing Algorithms, fill=blue!10 [ Randomization Partition: E2LSH [261], RPTree [262, 263], etc. , leaf, text width=27em ] [ Learned Partition: SPANN [264], etc. , leaf, text width=16em ] [ Navigable Partition: NSW [265], HNSW [266], etc. , leaf, text width=22em ] ] [Hardware-aware Indexing, fill=blue!10 [ DRAM-SSD: DiskANN [267], DiskANN++ [268], etc. , leaf, text width=23em ] [ Co-design: CXL-ANNS [269], FANNS [270], etc. , leaf, text width=21em ] ] ] ] ]
@@ 5.1.1 Model Compression
Pruning reduces the model size and computations by removing less important connections in the network. Pruning is categorized into structured pruning and unstructured pruning. Structure pruning usually removes weights in regular patterns, such as a rectangle block in the matrix or an entire channel, while unstructured pruning doesn’t impose such constraints. Consequently, structured pruning (e.g., LLM-Pruner [221]) is more hardware-friendly but more difficult to maintain model accuracy. While traditional pruning approaches require costly retaining process to preserve model capability, recent works like SparseGPT [222] and Wanda [223] have explored to perform unstructured pruning in one-shot.
@@ 5.1.2 Inference Acceleration
Context Compression methods enhance the inference efficiency by reducing the length of the context, especially the KV cache. Co-quantization of weights and activations, including KV cache, is an intuitive approach to compress the KV cache, which has been discussed in Section 5.1.1. Besides quantization, context pruning removes less important tokens in the context to reduce the computational cost. The effectiveness of this method is based on the observation that tokens have different impacts on the final output, and removing less important tokens won’t cause significant degradation of the model’s capability [233, 280, 234]. A typical line of work is to compress the context at the prefill stage based on different importance of tokens [230, 231, 232]. However, these methods are one-shot and cannot prune the KV cache when the context length continuously grows during token generation. To address this, Dynamic Context Pruning [233] uses a learnable mechanism to continuously determine and drop uninformative tokens. While the learnable mechanism introduces a fine-tuning overhead, Zhang et al. [234], proposes a token eviction strategy that can be applied without fine-tuning.
@@
Kernel Optimization is another approach towards LLM inference acceleration. Optimization for small-batch or single-batch inference is especially important for edge scenarios including the locally-deployed Personal LLM Agents. Existing works have revealed that the attention calculation becomes a bottleneck when the sequence length is long, since the complexity of attention scales quadratically with the sequence length, while that of the FFN scales linearly. Therefore, efficient attention kernels including FlashAttention [235, 236] and FlashDecoding++ [237] have been proposed to improve the speed of long-text inference. Besides, some works reduce the computational complexity of attention from the algorithm aspect. For example, Linformer [283] achieves linear complexity for self-attention in the prefill phase.
@@ 5.1.3 Memory Reduction
Although the compression-based methods are demonstrated to be able to effectively reduce the memory requirement of LLM inference, the accuracy degradation caused by compression are not negligible in some cases. To address this, FlexGen [240] designs an offloading strategy to fully utilize GPU, CPU and disk, together with a zig-zag scheduling scheme to support high-throughput inference under constrained GPU memory. This approach is orthogonal to compression-based methods, and thus can be jointly used to further reduce GPU memory footprints.
@@ 5.1.4 Energy Optimization
The energy consumption is a critical factor that affects the real-world deployment of an intelligent personal agent. An energy-consuming agent not only increases the deployment cost and carbon footprint, but also hurts the quality of experience (QoE) due to increased temperature and potential thermal throttling. While the inference of the LLM involves costly computations and memory accesses, it is important to optimize the energy efficiency of LLM inference.
@@ 5.2 Efficient Customization
The Personal LLM Agents may need to serve different users, different tasks, and different scenarios with the same base LLM, which requires efficient customization for each situation. There are mainly two ways to customize the behaviors of LLMs, one is feeding the LLM with different contextual prompts for in-context learning, and another is tuning the LLM with domain-specific data. Therefore, the efficiency of customization is primarily determined by the context loading efficiency and LLM fine-tuning efficiency.
@@ 5.2.1 Context Loading Efficiency
Another way to boost context loading is to reduce the bandwidth consumption during context data transmission. In some cases, pruning or discarding some tokens inevitably hurts the LLMs’ performance and loading the KV cache necessitates high bandwidth cost. CacheGen [252] addresses the challenges posed by context loading and it leverages the distinct characteristics of KV features across both tokens and layers thus introduces a novel KV encoder design. This encoder proficiently compresses the KV cache into a compact bitstream, effectively curtailing bandwidth demands while simultaneously reducing processing latency.
@@ 5.2.2 Fine-tuning Efficiency
Parameter-efficient fine-tuning (PEFT). A huge amount of parameters in LLMs make it costly to conduct full-parameter fine-tuning. Lots of efforts on parameter-efficient fine-tuning emerged to reduce LLMs’ training overhead. The fundamental concept of PEFT is to freeze the majority of parameters, focusing solely on training a limited set or introducing an adapter with significantly fewer parameters. A common practice is to introduce some adapters, i.e., small neural networks modules, into the existing network structure, including tuning hidden states [244, 245, 288], adding full layers [244] and prepending some prefix vectors into transformer architecture [289, 290, 291]. Liu et al. [292] also incorporates trainable vectors at the input layer, the performance of which highly depends on the capabilities of the underlying models. Some of these works fail to avoid extra adapter computation and introduce inference latency. LoRA [246] freezes all the model weights and augments each transformer layer with additional rank decomposition matrices, greatly reducing the memory and storage usage during fine-tuning without any additional inference latency. Another advantage of LoRA is that users can easily switch between different downstream tasks by simply adding or subtracting adapter matrices.
@@
Efficient Optimizer Design. Efficient optimizer design is another group of training/fine-tuning strategies which aims to accelerate the training or reduce the memory overhead during training. Sophia [248], a lightweight second-order optimizer, addresses the high cost and time required for LLM pre-training by providing a more efficient optimization process compared to commonly used methods like Adam and its variants. On the other hand, the huge number of parameters necessitates storing more activation and optimizer states especially in larger batch size, which places substantial memory demands. LOMO [247] presents a detailed analysis of the memory profile, throughput, and downstream performance of the proposed optimizer compared to other methods, demonstrating significant reductions in memory usage while maintaining training efficiency.
@@
Training Data Curation. Aforementioned approaches primarily focus on the process of training LLMs, while there are also some studies that aim to enhance the LLMs’ training performance from a distinct perspective, i.e., the amount and quality of training data. It has been demonstrated in phi-1 [249] that training the LLMs with a small amount of high-quality data can lead to significantly reduced training cost and achieve capabilities comparable to large-scale datasets and models. This challenges the traditional scaling laws in deep learning that emphasize larger datasets and models. Furthermore, phi-1.5 [250] and phi-2 [251] extend their focus on many other kinds of tasks such as common sense reasoning and language understanding, achieving comparable performance to models 5x and 25x larger, respectively.

### v2

@@ 4.2.2 Sensing Targets
  • Scene sensing is often readily perceptible but hold significant importance, leading to variations both in behavior and emphasis. For behavior instance, detecting a user in a library prompts the agent to adjust the phone to silent mode, while in a bar increasing the volume and activating vibration may be necessary. Similarly to emphasis, when a user is in a meeting room, the agent should focus more on tasks related to meeting content recording and work organization, whereas in a gym, emphasis should shift towards fitness plans and heart rate analysis. Previous work in scene awareness has employed various techniques [173], such as location-based approaches [174], audio or video analysis [175, 176], and sensor capabilities analyzing aspects like airflow through smartphone microphones to assess ventilation [140], or scene recognition achieved by analyzing macro photographs taken with the smartphone camera when placed near a surface [141]. Zhang et al. [147] let LLM understand 3D scenes through LLM-guided multiple viewpoint selection.
  • Occasion perception is more elusive in perception, and their impacts are relatively discreet. Earlier studies have identified differences in behavior and emotion recognition tasks across countries [177] and regions [178]. The national, ethnic, religious, and cultural backgrounds implied by the current user and setting are crucial. Perceiving others and objects in the current environment is equally vital. For example, previous work detected social scenarios based on sensor data, analyzing the behavior of socially anxious individuals in different social settings [179]. Other research delved into analyzing drinking-related social scenes using multiple sensors, even predicting the size and gender composition of drinking groups [180]. Additionally, studies explored the relations between sensor data, dietary habits, and social settings, revealing a strong association between binge eating and social environments, making it predictable [181]. Liang et al. [182] use LLM forecasting pedestrian flow through the analysis of public events.
@@
  • Short-term user sensing encompasses various aspects, including users’ routine actions [183], or specialized activities such as tooth brushing effectiveness [184], Ji et al. [145] found that even directly feeding IMU data to LLM can perform Human Activity Recognition (HAR) tasks. User states such as working or resting [160, 157], user health conditions [185, 139, 186], as well as user emotions [187, 156] and stress levels [188]. Recently, numerous studies have attempted to explore the applications of LLMs in the field of health monitoring [189, 190, 191]. Short-term sensing typically involve rapidly changing and shallow-level state information. Efficiently capturing such information can significantly enhance the context awareness of Personal LLM Agents.
  • Long-term user sensing mainly focus on the analysis of users’ profile and personality. Various approaches have been proposed to understand users’ work, study, and daily life. For instance, a study utilized sensor data from new smartphones to detect the prolonged psychological states of freshmen [192]. Another study demonstrated the capability to predict learning performance and social activities based on perception data [193]. Gao et al. [194] delve into the techniques to predict personality based on the intensity of physical activities. There is also research examining the relationship between sensor data and user career advancement [195], as well as a study that predicts user life satisfaction [196]. Furthermore, specific states of users have been a focus, including studies on the perception of mental illnesses [197, 198], such as one that predicts and analyzes schizophrenia [199], depression [190], and another that detects habits like smoking [200]. Lifelo et al. [191] utilized LLM to conduct psychological disorder analysis for a highly rare African language. Additionally, Ouyang and Srivastava [201] attempt to extract higher-level perceptual information from simple data. Long-term sensing involve deep and abstract information, containing the profound logic behind user behavior. These pieces of information are often more subtle, making perception and maintenance challenging. However, they constitute an essential aspect for advanced personal agents.
In terms of user sensing, there are also several LLM-based initiatives, such as employing LLM for recommendation tasks [202, 203], sentiment analysis with LLM [204], and the development of a personal doctor equipped with inquiry and perception capabilities [205].
@@ 5 Efficiency
Anagnostidis et al. [263], Zhang et al. [264], Ge et al. [265], etc. , leaf, text width=27em ] ] [Kernel Optimization, fill=blue!10 [ FlashAttention [266, 267], FlashDecoding++ [268], etc. , leaf, text width=24em ] ] [Speculative Decoding, fill=blue!10 [ Chen et al. [269], Leviathan et al. [270] , etc. , leaf, text width=20em ] ] ] [Memory Reduction (§5.1.3), fill=blue!10 [KV Quantization, fill=blue!10 [ ZeroQuant [249], SmoothQuant [250], etc. , leaf, text width=18em ] ] [KV Pruning, fill=blue!10 [ Anagnostidis et al. [263], Zhang et al. [264], etc. , leaf, text width=22em ] ] [Offloading, fill=blue!10 [ FlexGen [271], PowerInfer [272], Alizadeh et al. [273], etc. , leaf, text width=25em ] ] ] [Energy Optimization (§5.1.4), fill=blue!10 [Software Approaches, fill=blue!10 [ Same above, leaf, text width=6em ] ] [Hardware Approaches, fill=blue!10 [ NPU [274], TPU [275], FPGA [276], etc. , leaf, text width=18em ] ] ] ] [Efficient
@@
Manipulation (§5.3), fill=blue!10 [Search Efficiency (§5.3.1), fill=blue!10 [Indexing, fill=blue!10 [ Typical: Randomization Partition [286, 287], Learned
Partition [288], Navigable Partition [289], etc. , leaf, text width=23em ] [ Hardware-aware: DiskANN [290], CXL-ANNS [291], FANNS [292], etc. , leaf, text width=31em ] ] [Searching, fill=blue!10 [ Search Plan [293, 294, 295, 296], Metadata Filtering [295, 297], etc. , leaf, text width=29em ] [ Execution: GPU [298, 296], SIMD [298, 296, 299],
OPENMP [298, 296], Distributed [300, 293], etc. , leaf, text width=22em ] ] ] [Workflow Efficiency (§5.3.2), fill=blue!10 [Pipelining, fill=blue!10 [ RaLMSpec [301], PipeRAG [302], etc. , leaf, text width=16em ] ] [Caching, fill=blue!10 [ RAGCache [303], GRITLM [304], etc. , leaf, text width=16em ] ] ] ] ]
@@ 5.1.1 Model Compression
While integer quantization methods such as INT4 and INT8 remain mainstream solutions in current deployment practice, there has been a new trend of low-bit floating point quantization, such as FP4 and FP8. One reason is that floating point quantization can achieve comparable or even higher accuracy than integer quantization [312, 313, 314]. Besides, floating point quantization is possible to achieve higher computational performance on both cloud GPUs like NVIDIA H100 with dedicated computing support, and mobile GPUs [315].
Pruning reduces the model size and computations by removing less important connections in the network. Pruning is categorized into structured pruning and unstructured pruning. Structure pruning usually removes weights in regular patterns, such as a rectangle block in the matrix or an entire channel, while unstructured pruning doesn’t impose such constraints. Consequently, structured pruning (e.g., LLM-Pruner [251]) is more hardware-friendly but more difficult to maintain model accuracy. While traditional pruning approaches require costly retaining process to preserve model capability, recent works like SparseGPT [252] and Wanda [253] have explored to perform unstructured or semi-structured pruning in one-shot.
@@ 5.1.2 Inference Acceleration
Context Compression methods enhance the inference efficiency by reducing the length of the context, especially the KV cache. Co-quantization of weights and activations, including KV cache, is an intuitive approach to compress the KV cache, which has been discussed in Section 5.1.1. Besides quantization, context pruning removes less important tokens in the context to reduce the computational cost. The effectiveness of this method is based on the observation that tokens have different impacts on the final output, and removing less important tokens won’t cause significant degradation of the model’s capability [263, 318, 264, 265]. A typical line of work is to compress the context at the prefill stage based on different importance of tokens [260, 261, 262]. However, these methods are one-shot and cannot prune the KV cache when the context length continuously grows during token generation. To address this, Dynamic Context Pruning [263] uses a learnable mechanism to continuously determine and drop uninformative tokens. While the learnable mechanism introduces a fine-tuning overhead, Zhang et al. [264], proposes a token eviction strategy that can be applied without fine-tuning.
@@
Kernel Optimization is another approach towards LLM inference acceleration. Optimization for small-batch or single-batch inference is especially important for edge scenarios including the locally-deployed Personal LLM Agents. Existing works have revealed that the attention calculation becomes a bottleneck when the sequence length is long, since the complexity of attention scales quadratically with the sequence length, while that of the FFN scales linearly. Therefore, efficient attention kernels including FlashAttention [266, 267] and FlashDecoding++ [268] have been proposed to improve the speed of long-text inference. Some works also reduce the computational complexity of attention from the algorithm aspect. For example, Linformer [321] achieves linear complexity for self-attention in the prefill phase. Besides, reducing dequantization overhead also provides significant performance improvement as demonstrated by LUT-GEMM [322].
@@ 5.1.3 Memory Reduction
Although the compression-based methods are demonstrated to be able to effectively reduce the memory requirement of LLM inference, the accuracy degradation caused by compression are not negligible in some cases. To address this, FlexGen [271] designs an offloading strategy to fully utilize GPU, CPU and disk, together with a zig-zag scheduling scheme to support high-throughput inference under constrained GPU memory. This approach is orthogonal to compression-based methods, and thus can be jointly used to further reduce GPU memory footprints. Another line of work, including PowerInfer [272] and Alizadeh et al. [273], reduces swapping overhead in low-batch inference by predicting contextual sparsity as inspired in [326].
@@ 5.1.4 Energy Optimization
The energy consumption is a critical factor that affects the real-world deployment of LLM agents given LLM’s costly computations and memory accesses. An energy-consuming agent not only increases the runtime cost and carbon footprint, but also hurts the quality of experience (QoE) due to increased temperature and shorten battery lifespan. Therefore, it is important to optimize the energy efficiency of LLM inference.
@@
Yet, the research on energy efficiency of LLM inference is still far from insufficient due to the complexity of hardware deployment and the volatility of energy measurement and analysis. There have be several studies that focus on this topic, such as evaluaing LLMs’ inference energy on GPUs [328, 329], edge devices [330] and carbon footprint of LLMs in datacenters [331]. Other works tent to present fast energy prediction method for LLM inference, such as IrEne [332], which conducted layer-level energy analysis on Transformer-based NLP models and gave an interpretable and extensible energy prediction system. However these prediction models are only for GPU host backends and lack of generalization to other hardware platforms such as mobile phones where Personal LLM Agents are more likely to be deployed.
@@ 5.2 Efficient Customization
The Personal LLM Agents may need to serve different users, different tasks, and different scenarios with the same base LLM, which requires efficient customization for each situation. There are mainly two ways to customize the behaviors of LLMs; one is feeding the LLM with different contextual prompts for in-context learning, and another is tuning the LLM with domain-specific data. Therefore, the efficiency of customization is primarily determined by the context loading efficiency and LLM fine-tuning efficiency.
@@ 5.2.1 Context Loading Efficiency
Another way to boost context loading is to reduce the bandwidth consumption during context data transmission. In some cases, pruning or discarding some tokens inevitably hurts the LLMs’ performance and loading the KV cache necessitates high bandwidth cost. CacheGen [285] addresses the challenges posed by context loading and it leverages the distinct characteristics of KV features across both tokens and layers thus introduces a novel KV encoder design. This encoder proficiently compresses the KV cache into a compact bitstream, effectively curtailing bandwidth demands while simultaneously reducing processing latency. Besides, given the fact that different input prompts may have overlapping text segments, Gim et al. [333] proposes Prompt Cache to reuse attention states across prompts. By pre-computing and storing the attention states of frequently occurring text, the framework can efficiently reuse them when these segments appear in new prompts, thus accelerating the inference process.
@@ 5.2.2 Fine-tuning Efficiency
Parameter-efficient fine-tuning (PEFT). A huge amount of parameters in LLMs make it costly to conduct full-parameter fine-tuning. Lots of efforts on parameter-efficient fine-tuning emerged to reduce LLMs’ training overhead. The fundamental concept of PEFT is to freeze the majority of parameters, focusing solely on training a limited set or introducing an adapter with significantly fewer parameters. A common practice is to introduce some adapters, i.e., small neural networks modules, into the existing network structure, including tuning hidden states [277, 278, 334], adding full layers [277] and prepending some prefix vectors into transformer architecture [335, 336, 337]. Liu et al. [338] also incorporates trainable vectors at the input layer, the performance of which highly depends on the capabilities of the underlying models. Some of these works fail to avoid extra adapter computation and introduce inference latency. LoRA [279] freezes all the model weights and augments each transformer layer with additional rank decomposition matrices, greatly reducing the memory and storage usage during fine-tuning without any additional inference latency. Another advantage of LoRA is that users can easily switch between different downstream tasks by simply adding or subtracting adapter matrices. $\mathtt{(IA)^{3}}$ [339] explores element-wise multiplication of the model’s activations against learned vectors. It introduces learned vectors which rescale the keys and values in attention mechanisms, and the inner activations in position-wise feed-forward networks. By only training the vectors, $\mathtt{(IA)^{3}}$ could maintain the performance with much less computation.
@@
Efficient Optimizer Design. Efficient optimizer design is another group of training/fine-tuning strategies which aims to accelerate the training or reduce the memory overhead during training. Sophia [281], a lightweight second-order optimizer, addresses the high cost and time required for LLM pre-training by providing a more efficient optimization process compared to commonly used methods like Adam and its variants. On the other hand, the huge number of parameters necessitates storing more activation and optimizer states especially in larger batch size, which places substantial memory demands. LOMO [280] presents a detailed analysis of the memory profile, throughput, and downstream performance of the proposed optimizer compared to other methods, demonstrating significant reductions in memory usage while maintaining training efficiency. Zhao et al. [340] propose HiZOO, aimed at leveraging the diagonal Hessian to enhance zeroth-order optimizer for fine-tuning LLMs. It avoids the expensive memory cost with one more forward pass per step.
@@
Training Data Curation. Aforementioned approaches primarily focus on the process of training LLMs, while there are also some studies that aim to enhance the LLMs’ training performance from a distinct perspective, i.e., the amount and quality of training data. It has been demonstrated in phi-1 [282] that training the LLMs with a small amount of high-quality data can lead to significantly reduced training cost and achieve capabilities comparable to large-scale datasets and models. This challenges the traditional scaling laws in deep learning that emphasize larger datasets and models. Furthermore, phi-1.5 [283] and phi-2 [284] extend their focus on many other kinds of tasks such as common sense reasoning and language understanding, achieving comparable performance to models 5x and 25x larger, respectively. Similarly, TinyGSM [341] introduced a synthesized dataset with a few amount (12.3M) of samples on grade school math, which led to remarkable accuracy when tuning small language models with the dataset.